OpenStudy (anonymous):

Help will medal and fan Write an expression to represent the total volume of two different sized boxes as a sum of cubes if one of the boxes has sides with a length of 1 foot and the other has sides with a length of x feet.

OpenStudy (studygurl14):

First off, what is the formula for the volume of a cube?

OpenStudy (anonymous):

I don't know

OpenStudy (jordanloveangel):

The total volume is x^3 + 1^3 = x^3 + 1 x^3 + 1 = (x + 1)(x^2 - x + 1) Let x = 3 x^3 + 1 = 3^3 + 1 = 28 You can confirm this by using the factors. (3 + 1) ( 3^2 - 3 + 1) (4 ) (9 - 3 + 1) 4 * 7 = 28 i think

OpenStudy (studygurl14):

The formula is the side length of the cube to the power of three, aka cubed. So if the side length is 1, what is the volume?

OpenStudy (anonymous):

1

OpenStudy (jordanloveangel):

thx